### v1

# SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking

Abstract
The rise of misinformation, exacerbated by Large Language Models (LLMs) like GPT and Gemini, demands robust fact-checking solutions, especially for low-resource languages like Vietnamese. Existing methods struggle with semantic ambiguity, homonyms, and complex linguistic structures, often trading accuracy for efficiency. We introduce SemViQA, a novel Vietnamese fact-checking framework integrating Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC). Our approach balances precision and speed, achieving state-of-the-art results with 78.97% strict accuracy on ISE-DSC01 and 80.82% on ViWikiFC, securing 1st place in the UIT Data Science Challenge. Additionally, SemViQA Faster improves inference speed 7× while maintaining competitive accuracy. SemViQA sets a new benchmark for Vietnamese fact verification, advancing the fight against misinformation. The source code is available at: https://github.com/DAVID-NGUYEN-S16/SemViQA.
SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking
**footnotetext: Equal contribution.footnotetext: Preprint
1 Introduction
The rapid advancement of large language models (LLMs), such as OpenAI’s ChatGPT, Google Gemini Team et al. (2024), Llama3.1 Touvron et al. (2023), Qwen2.5 Qwen et al. (2025), DeepSeek V3, DeepSeek-AI et al. (2024), Phi3.5 Abdin et al. (2024) has significantly improved information retrieval and processing across various domains. However, a major challenge with these systems is their tendency to generate factually incorrect or hallucinated content seemingly plausible information that lacks factual grounding  Soleimani et al. (2020). This issue is particularly critical in domains requiring high accuracy, such as healthcare, law, and journalism, where misinformation can have serious consequences. Consequently, developing reliable fact-checking systems capable of retrieving and evaluating evidence from real-world sources has become an urgent need in Natural Language Processing (NLP).
Although fact verification has been extensively studied in high-resource languages like English, with a primary focus on datasets such as FEVER Thorne et al. (2018), LIAR Wang (2017), PubHealthTab Akhtar et al. (2022), and TabFact Chen et al. (2020), extending these methods to low-resource languages like Vietnamese remains a significant challenge. Transformer-based approaches, including BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), have shown strong performance in fact verification, yet their adaptation to Vietnamese is still underexplored. For Vietnamese fact verification, ViNSV Tran et al. (2024) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval. However, SBERT’s 256-token input limitation leads to information loss in long-context claims, making it particularly ineffective for complex cases requiring multi-sentence reasoning or cross-referencing distant evidence. As a result, the system struggles with hard samples that demand deeper contextual understanding, ultimately reducing fact verification accuracy. In contrast, graph-based reasoning methods in English Zhong et al. (2020) have shown notable potential for semantic inference but are often hindered by high computational costs, restricting their real-world applicability. Moreover, traditional retrieval techniques like TF-IDF and BM25, though widely used, rely heavily on token frequency and exact keyword matching, making them less effective at capturing the nuanced semantics of complex claims. While some recent work employing large language models (LLMs) (Huo et al., 2023; Schimanski et al., 2024) has shown promise in addressing hard retrieval questions, these approaches typically require considerably more computing resources and longer processing times compared to conventional methods. Thus, a trade-off persists: models offering high accuracy tend to operate slowly, whereas faster models often compromise on accuracy.
To address the challenges in fact verification, we introduce SemViQA, a system designed to optimize both accuracy and efficiency. It comes in two variants: SemViQA Standard, which prioritizes accuracy, and SemViQA Faster, which significantly improves inference speed while maintaining strong performance. The system is built upon three core components:
  1. 1.
    Data Processing: A specialized preprocessing pipeline designed to efficiently handle long-token sequences, a known limitation of Transformer-based models. Our approach preserves the semantic integrity of input claims while ensuring optimal tokenization for downstream processing.
  2. 2.
    Semantic-based Evidence Retrieval (SER): A hybrid retrieval framework combining TF-IDF for efficient keyword-based matching and a Question Answering Token Classifier (QATC) for semantic-based evidence selection. This approach balances speed and accuracy by prioritizing TF-IDF for common cases while activating QATC selectively for hard samples, significantly improving precision while maintaining computational efficiency.
  3. 3.
    Two-step Verdict Classification (TVC): A hierarchical classification strategy leveraging Focal Loss Lin et al. (2018) and Cross-Entropy Loss to enhance claim verification accuracy. This approach first determines whether a claim is Supported, Refuted, or Not Enough Information (NEI) and subsequently refines the decision boundary for hard-to-classify cases, reducing misclassification and improving robustness.
We evaluate SemViQA on two of the largest Vietnamese fact verification datasets, ISE-DSC01 and ViWikiFC Le et al. (2024). Our system achieves 78.97% strict accuracy on ISE-DSC01  [^1] https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC, outperforming existing baselines. These results highlight SemViQA’s potential to improve the accuracy of fact verification in Vietnamese, contributing to misinformation mitigation and enhancing transparency in digital content.
The remainder of this paper is structured as follows: Section 2 discusses related work. Section 3 details our proposed methodology, including SER, TVC, and QATC. Section 4 presents experimental results and performance evaluation. Finally, Section 5 concludes the paper and outlines future research directions.
Figure 1: SemViQA: A Three-Stage Method for semantic-based evidence retrieval and two-step verdict classification, where $P_{2}$ and $P_{3}$ represent the probabilities of the two-class and three-class classifications, respectively, and $\hat{y}_{\text{2}}$ and $\hat{y}_{\text{3}}$ denote their corresponding predictions.
2 Related Works
Advancements in Natural Language Processing (NLP) have significantly improved information verification and evidence extraction. Early methods, such as the Neural Semantic Matching Network (NSMN) Nie et al. (2018), leveraged BiLSTM Graves and Schmidhuber (2005) and WordNet-based features to enhance verification accuracy but struggled with complex sentence relations due to LSTM’s sequential nature.
The introduction of transformer models, notably BERT Devlin et al. (2019), transformed claim verification by enabling bidirectional contextual encoding Soleimani et al. (2020); Zhou et al. (2019); Malon (2018). While achieving state-of-the-art results on FEVER Aly et al. (2021); Lin et al. (2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), BERT’s input length limitation makes it less effective for long-context fact-checking.
Graph-based models have further advanced multi-step reasoning. Approaches such as GCN-based reasoning Zhong et al. (2020); Thorne et al. (2018) and AdMIRaL’s logic-driven document retrieval Aly and Vlachos (2022) improve evidence sufficiency but demand high computational resources.
For Vietnamese fact verification, research remains limited. ViNSV Tran et al. (2024) combined BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but struggled with complex reasoning due to static embeddings.
Building on prior work, SemViQA optimizes fact verification by integrating fast keyword-based retrieval (TF-IDF) with semantic reasoning (QATC) and a hierarchical classification strategy, ensuring high accuracy, fast inference, and scalability for real-world applications.
3 SemViQA - Semantic Vietnamese Question Answering
Our proposed method, SemViQA, employs a three-stage pipeline to enhance fact verification, as illustrated in Figure 1. The process begins with a preprocessing stage, where input data is restructured to ensure compatibility with retrieval and classification models. Subsequently, evidence retrieval is performed using TF-IDF for straightforward cases, while the Question Answering Token Classifier (QATC) refines evidence selection for more complex claims. Finally, in the claim classification stage, a hierarchical approach integrates Cross-Entropy Loss and Focal Loss  Lin et al. (2018) to improve both accuracy and robustness.
3.1 Data Processing
Figure 2: Graph representing the lengths of contexts.
Before implementing data processing solutions, we conducted a thorough analysis of the input data to understand its characteristics and challenges. One of the prominent issues we encountered is the token limit of current Vietnamese BERT models, as clearly illustrated in Figure 2.
During data processing, we observed that the input context has a considerable length, with the longest context extending up to 4803 tokens. This presents a challenge for data processing given the limitations of 256 tokens for classification and 512 tokens for the Question Answering model.
[…]
Figure 3: Long context processing solution.
From a long context, we proceed to split it into smaller segments. We then slide through each segment and check the number of tokens in each segment. If the token count does not exceed 400, we append the segment to the subcontext. When the subcontext has a token count greater than 400 (see Figure 3), we check if the evidence sentence is contained within the subcontext. If it is, then the subcontext is considered to contain that evidence. Otherwise, if it is not present, the evidence for that subcontext is assigned as an empty string. This process allows us to create new data samples with contexts containing fewer than 512 tokens, which is beneficial for training the model more effectively. This ensures that no segment is omitted, unlike stride-based cutting methods, which may result in evidence sentences being split between segments within a complete paragraph.
3.2 Semantic-based Evidence Retrieval (SER)
In addressing this problem, focusing on evidence retrieval is a crucial and decisive part. Accurate classification can only be performed when the evidence is correct. After analyzing and training on the data, we found that methods like TF-IDF and BM25 are effective for simpler samples but struggle with more complex samples.
Therefore, we have divided the evidence retrieval process into two parts:
Part 1: Using TF-IDF: We begin by segmenting the context into smaller passages and pairing each segment with the claim to generate structured data samples, which then undergo a preprocessing pipeline including noise removal and tokenization using ViTokenizer  [^2] https://github.com/trungtv/pyvi to standardize the input. A key observation is that TF-IDF performs well on simple claims, particularly those classified as refuted, but struggles with claims requiring deeper reasoning, as it relies solely on keyword matching without capturing semantic relationships. To address this limitation, we introduce an enhanced strategy to enrich contextual information, where segments containing fewer than 60% of the claim’s token count are merged with the preceding segment to supplement missing information. This is particularly beneficial for short or incomplete sentences where crucial words may be absent from the claim, leading to improved evidence retrieval accuracy. After retrieving potential evidence using TF-IDF, we rank the retrieved segments and apply a Confidence Threshold to classify them into easy and challenging cases, as detailed in Appendix C. While simple cases can be resolved directly using TF-IDF, more complex cases are forwarded to Part 2 for further analysis, ensuring higher classification accuracy.
Part 2: Using the Question Answering Token Classifier Model: The subcontext is also used for model QA inference instead of inferring the full context. In this process, three situations can occur: The first, when multiple subcontexts generate different responses, we interpret this as the model exhibiting confusion in handling evidence responses. In this case, we collect all retrieved evidence and apply TF-IDF to select the most appropriate one. The second, when only one piece of evidence appears across multiple subcontexts, we interpret this as the model being highly confident in its prediction. The third, when the model does not retrieve any evidence, we use TF-IDF to extract relevant evidence as input for our TVC method. By combining these methods, we have achieved the best results in accurately and reliably retrieving evidence.
3.3 Two-step Verdict Classification (TVC)
Our input data consists of claim-evidence pairs, where the claim is the statement that needs to be verified, and the evidence comprises information used to support (supported), refute (refuted), or indicate insufficient information (NEI) regarding the claim. To address this problem, we undertake two main tasks as follows:
Three-Class Classification: The first step involves categorizing each claim into one of three categories: Supported, Refuted, or Not Enough Information (NEI). Given a claim-evidence pair $(C,E)$, we employ a classification function:
$P_{\text{3}}=f_{\text{Three-Class}}(C,E),$ (1)
where
$P_{\text{3}}=\left[P_{\text{SUP}},P_{\text{REF}},P_{\text{NEI}}\right]$
represents the predicted probabilities. Here, $f_{\text{Three-Class}}$ denotes a neural network model based on an advanced BERT-based architecture, trained to classify each claim into one of the three categories. The model leverages a fine-tuned BERT-derived approach and is optimized using Cross Entropy Loss, ensuring that the predicted probability distribution closely aligns with the true labels.
Binary Classification: If the three-class model does not predict NEI, an additional classification step is applied to further refine the decision between Supported and Refuted. In this step, we use another classification function:
$P_{\text{2}}=f_{\text{Binary-Class}}(C,E),$ (2)
where
$P_{\text{2}}=\left[P_{\text{SUP}},P_{\text{REF}}\right].$
Here, $f_{\text{Binary-Class}}$ is a neural network based on an advanced BERT-based architecture, specifically trained to distinguish between Supported and Refuted claims. Unlike the three-class model, this binary classifier leverages a fine-tuned BERT-derived model and is trained using Focal Loss  Lin et al. (2018) to mitigate class imbalance and improve performance on challenging examples.
Finally, the predicted class $\hat{y}$ is determined based on the following decision rule:
$\hat{y}=\begin{cases}\text{NEI},&\text{if }\hat{y}_{\text{3}}=\text{NEI}\\ \hat{y}_{\text{2}},&\text{if }P_{\text{2}}>P_{\text{3}}\text{ and }\hat{y}_{\text{3}}\neq\hat{y}_{\text{2}}\\ \hat{y}_{\text{3}},&\text{if }P_{\text{3}}\geq P_{\text{2}}\text{ or }\hat{y}_{\text{3}}=\hat{y}_{\text{2}}.\end{cases}$ (3)
This method effectively balances the strengths of both classification models, ensuring improved accuracy for complex claims while maintaining high performance on straightforward cases.
3.4 Question Answering Token Classifier (QATC)
Typical QA models are trained only to predict the start and end positions of an answer. Here, we extend this by introducing a Token Classification task, where we assign a label of 1 to tokens within the answer and 0 otherwise. This enhances the model’s ability to identify relevant tokens for answer extraction.
Procedure: We input the context and claim into the model, extract token features, and apply two linear layers to predict answer start and end probabilities. The default QA loss is the Cross Entropy loss:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}t_{i}\log(p_{i}),$ (4)
where $N$ is the number of classes, $t_{i}$ is the true label, and $p_{i}$ is the softmax probability for the $i^{\text{th}}$ class.
3.4.1 Token Classification Task
Inspired by Rationale Tagging Ju et al. (2019), we predict whether each token is part of the answer (label 1) or not (label 0). For NEI (Not Enough Information) cases, all tokens are labeled 0. To achieve this, we introduce an additional fully connected (FC) layer that computes classification probabilities:
$p_{t}^{r}=\sigma\bigl{(}w_{2}\cdot\text{ReLU}(W_{1}h_{t})\bigr{)},$ (5)
where $h_{t}$ is the output representation of token $t$ from RoBERTa Zhuang et al. (2021); Aly et al. (2021). The loss for this task is the Binary Cross Entropy loss:
$\mathcal{L}_{RT}=-\frac{1}{T}\sum_{t=1}^{T}\bigl{[}y_{t}^{r}\log(p_{t}^{r})+(1-y_{t}^{r})\log(1-p_{t}^{r})\bigr{]}.$ (6)
3.4.2 Rationale Regularization Loss
To encourage sparsity and continuity in rationale predictions, we introduce a Rationale Regularization Loss $\mathcal{L}_{RR}$, which consists of:
1. Sparsity Loss: Encourages the model to minimize the number of tokens selected as rationales:
$\mathcal{L}_{sparse}=\frac{1}{T}\sum_{t=1}^{T}m_{t}p_{t}^{r},$ (7)
where $m_{t}$ is a mask ensuring that the loss is computed only on non-padding tokens.
2. Continuity Loss: Encourages smoother rationale selection by minimizing sudden jumps between selected tokens:
$\mathcal{L}_{cont}=\frac{1}{T-1}\sum_{t=1}^{T-1}m_{t}m_{t+1}(p_{t}^{r}-p_{t+1}^{r})^{2}.$ (8)
The total rationale regularization loss is:
$\mathcal{L}_{RR}=\lambda_{sparse}\mathcal{L}_{sparse}+\lambda_{cont}\mathcal{L}_{cont}.$ (9)
3.4.3 Combined Loss Function
The final objective function integrates all components:
$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{RR}+\alpha\mathcal{L}_{RT},$ (10)
where $\alpha$ controls the weight of the rationale tagging loss, and $\lambda_{sparse}$, $\lambda_{cont}$ are hyperparameters for regularization.
4 Experiments
4.1 Dataset
ISE-DSC01, used in the UIT Challenge 2023, is designed for AI evaluation in information extraction and verification. It comprises training, public test, and private test sets with diverse claim-context pairs, posing challenges in evidence retrieval and fact-checking.
ViWikiFC Le et al. (2024) was developed to combat misinformation, featuring over 20,000 claims from Wikipedia. Unlike ISE-DSC01, it explicitly includes evidence for the “nei” (Not Enough Information) label, enhancing real-world applicability.
ISE-DSC01 ViWikiFC
Train 37,967 16,738
Dev 4,794 2,090
Test 5,396 2,091
Table 1: Overview of the datasets used in our experiments
Table 1 summarizes the datasets. For ViWikiFC, experiments were conducted on all three sets: training, public test, and private test.
Method Parameter ViWikiFC ISE-DSC01 Avg Strict Acc
ER VC Strict Acc VC Acc ER Acc Time (s) Strict Acc VC Acc ER Acc Time (s)
TF-IDF InfoXLMlarge 560M 75.56 82.21 90.15 131 73.59 78.08 76.61 378 74.58
XLM-Rlarge 560M 76.47 82.78 90.15 134 75.61 80.50 78.58 366 76.04
Ernie-Mlarge 560M 75.56 81.83 90.15 144 78.19 81.69 80.65 403 76.88
BM25 InfoXLMlarge 560M 70.44 79.01 83.50 130 72.09 77.37 75.04 320 71.27
XLM-Rlarge 560M 70.97 78.91 83.50 132 73.94 79.37 76.95 333 72.46
Ernie-Mlarge 560M 70.21 78.29 83.50 141 76.58 80.76 79.02 381 73.40
SBert InfoXLMlarge 838M 74.99 81.59 89.72 195 71.20 76.59 74.15 915 73.10
XLM-Rlarge 838M 75.80 82.35 89.72 194 72.85 78.78 75.89 835 74.33
Ernie-Mlarge 838M 75.13 81.44 89.72 203 75.46 79.89 77.91 920 75.30
QA-based approaches VC
ViMRClarge InfoXLMlarge 1120M 77.28 81.97 92.49 3778 54.36 64.14 56.84 9798 65.82
XLM-Rlarge 1120M 78.29 82.83 92.49 3824 53.98 66.70 57.77 9809 66.14
Ernie-Mlarge 1120M 77.38 81.92 92.49 3785 56.62 62.19 58.91 9833 67.00
InfoXLMlarge InfoXLMlarge 1120M 78.14 82.07 93.45 4092 53.50 63.83 56.17 10057 65.82
XLM-Rlarge 1120M 79.20 83.07 93.45 4096 53.32 66.70 57.25 10066 66.26
Ernie-Mlarge 1120M 78.24 82.21 93.45 4102 56.34 62.36 58.69 10078 67.29
LLM
Qwen2.5-1.5B-Instruct 1.5B 51.03 65.18 78.96 7665 59.23 66.68 65.51 19780 55.13
Qwen2.5-3B-Instruct 3B 44.38 62.31 71.35 12123 60.87 66.92 66.10 31284 52.63
LLM VC
Qwen2.5-1.5B-Instruct InfoXLMlarge 2B 66.14 76.47 78.96 7788 64.40 68.37 66.49 19970 65.27
XLM-Rlarge 2B 67.67 78.10 78.96 7789 64.66 69.63 66.72 19976 66.17
Ernie-Mlarge 2B 66.52 76.52 78.96 7794 65.70 68.37 67.33 20003 66.11
Qwen2.5-3B-Instruct InfoXLMlarge 3.5B 59.88 72.50 71.35 12246 65.72 69.66 67.51 31477 62.80
XLM-Rlarge 3.5B 60.74 73.08 71.35 12246 66.12 70.44 67.83 31483 63.43
Ernie-Mlarge 3.5B 60.02 72.21 71.35 12251 67.48 70.77 68.75 31512 63.80
SER Faster (ours) TVC (ours)
TF-IDF + ViMRClarge Ernie-Mlarge 1680M 79.44 82.93 94.60 410 78.32 81.91 80.26 995 78.88
TF-IDF + InfoXLMlarge 1680M 79.77 83.07 95.03 487 78.37 81.91 80.32 925 79.07
SER (ours) TVC (ours)
TF-IDF + ViMRClarge InfoXLMlarge 1680M 80.25 83.84 94.69 2731 75.13 79.54 76.87 5191 77.69
XLM-Rlarge 1680M 80.34 83.64 94.69 2733 76.71 81.65 78.91 5219 78.53
Ernie-Mlarge 1680M 79.53 82.97 94.69 2733 78.97 82.54 80.91 5225 79.25
TF-IDF + InfoXLMlarge InfoXLMlarge 1680M 80.68 83.98 95.31 3860 75.13 79.60 76.87 5175 77.91
XLM-Rlarge 1680M 80.82 83.88 95.31 3843 76.74 81.71 78.95 5200 78.78
Ernie-Mlarge 1680M 80.06 83.17 95.31 3891 78.97 82.49 80.91 5297 79.52
Table 2: Performance comparison on the ViWikiFC test set and the ISE-DSC01 private-test dataset. The results highlight differences among models based on several criteria: Strict Accuracy (Strict Acc), Veracity Classification Accuracy (VC Acc), and Evidence Retrieval Accuracy (ER Acc). Time represents the total inference time required to generate the complete results. Parameter indicates the total number of parameters used in each task. The results highlighted in blue indicate that our SER Faster method achieves the highest performance among all methods, except for the standard SER method.
4.2 Experimental Setup
We optimized our model through extensive experiments on an NVIDIA A100 GPU, fine-tuning key hyperparameters while maintaining consistent settings across all runs. The final hyperparameters, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC. Details are provided in Appendix B. For evaluation, all methods were tested on an NVIDIA T4 instance in Kaggle to ensure fairness.
Additionally, the large language model was fine-tuned on a distributed setup with 4×H100 GPUs, utilizing a structured prompt reformulation process. The dataset, originally in raw text and labels, was transformed into a prompt-based format optimized for LLM fine-tuning. This ensured effective task-specific learning while aligning with pre-trained knowledge. Further details on training setup, prompt design, and preprocessing are in Appendix B.
4.3 Main Results
The results in Table 2 demonstrate that SemViQA outperforms previous methods in Vietnamese fact-checking tasks. Specifically, our model achieves the highest Strict Accuracy, reaching 80.82% on ViWikiFC and 78.97% on ISE-DSC01, establishing a new benchmark for automated fact-checking systems in Vietnamese language.
4.3.1 Performance Comparison
a) Handling Long Token Sequences in Fact-Checking
One major limitation of conventional Question Answering (QA) models in fact verification tasks is their inability to process long-context claims due to token length constraints. Transformer-based models such as ViMRClarge, InfoXLMlarge Chi et al. (2021), and XLM-Rlarge Conneau et al. (2020), Ernie-Mlarge Ouyang et al. (2021) are restricted to a maximum input length of 512 tokens. However, real-world fact-checking datasets, such as ISE-DSC01, often contain contexts exceeding 4800 tokens, causing a significant drop in performance for QA-based approaches due to their inability to handle long-token sequences. This limitation prevents these models from effectively leveraging the full evidence available, leading to incomplete or inaccurate fact verification.
To address this issue, our proposed SemViQA model incorporates a retrieval-based approach capable of handling long-token sequences efficiently, as illustrated in Figure 3. On ISE-DSC01, where long contexts are essential for verification, SemViQA significantly outperforms traditional QA models by effectively capturing and utilizing extended contextual information. This demonstrates that the long-token issue is a critical bottleneck in fact verification tasks.
In contrast, on the ViWikiFC dataset, where the token length is closer to 512 tokens, QA-based models achieve competitive performance, further confirming that the long-token issue is a major factor in performance degradation. However, when the long-token limitation does not exist on the ViWikiFC dataset, integrating our SER approach into a QA-based model proves to be an effective solution. By leveraging SER, we achieve approximately 1.86% higher ER Accuracy on ViWikiFC, demonstrating its efficiency in optimizing fact verification when token constraints are minimized.
These results highlight that long-token constraints are a fundamental challenge in fact verification. Our approach not only mitigates this issue through retrieval-based processing in SemViQA but also enhances QA models with QACT when long-token constraints are less severe, ensuring improved performance across diverse datasets.
b) Performance and Inference Time Optimization
One of the key advancements of SemViQA over other methods is its significant reduction in inference time without compromising accuracy. Key observations include:
  • The average inference time of SemViQA on ISE-DSC01 is 5200s, whereas large LLM-based models such as Qwen2.5-3B-Instruct Qwen et al. (2024) require over 31,000s, making SemViQA at least 6 times faster than these large language model approaches.
  • Compared to ViMRClarge [^3] https://huggingface.co/nguyenvulebinh/vi-mrc-large (9800s on ISE-DSC01), SemViQA reduces inference time by nearly 50% while maintaining higher performance in both Strict Accuracy and Veracity Classification Accuracy.
  • Methods based on SBERT Reimers and Gurevych (2019b) or BM25 have lower inference times but fail to maintain high accuracy, particularly when dealing with multi-step reasoning tasks. SemViQA achieves a better balance between classification performance and processing speed, ensuring practical deployment in real-world environments.
Figure 4: Comparison of method performance, balancing accuracy and inference time. Each retrieval method is evaluated based on its highest achieved score, while the total inference time across both datasets is reported to highlight efficiency. Further details can be found in Table 2.
SemViQA Faster: We introduce an optimized version of SemViQA that achieves up to 7× faster inference while maintaining high accuracy. As shown in Figure  4, the inference speed of SER-Faster closely matches that of traditional methods while still outperforming competing solutions in both efficiency and accuracy. To achieve this speed improvement, we process all subcontexts within a sample as a batch, enabling batch inference (see Section 3.1 for details on subcontexts). While this approach results in a slight performance drop compared to the standard SemViQA model, the trade-off is minimal considering the significant reduction in processing time. This efficiency enhancement makes SemViQA Faster highly suitable for real-world deployment, allowing seamless integration into large-scale fact-checking applications.
4.3.2 Analysis of Confidence Threshold in SemViQA
Figure 5: Impact of confidence threshold on evidence retrieval accuracy in SemViQA.
The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process. Analysis from Figure 5 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%). However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively. Setting an optimal threshold in the range of 0.4 – 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
4.3.3 Comparison with other results in the competition
Methods Strict Acc VC Acc ER Acc
SemViQA 78.97 82.54 80.91
DS@UIT Dynasty 78.05 84.76 80.13
URA_FNU 77.87 83.71 79.96
Plain Sailing 77.09 82.33 78.31
ViNSV 76.33 81.67 78.11
Table 3: Comparison of results with the top 5 teams in the competition
The results presented in Table 3 indicate that our SemViQA approach outperforms other competing teams, achieving the highest Strict Accuracy and demonstrating exceptional effectiveness in information processing and verification. This achievement highlights SemViQA’s capability to deliver significantly more accurate and reliable results.
5 Conclusion and Future Works
We introduced SemViQA, a Vietnamese fact-checking framework that integrates Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC) to enhance claim verification. Our approach outperforms existing methods, including LLMs, TF-IDF, BM25, SBERT, and QA-based models, particularly in handling long-token sequences and complex reasoning tasks. Extensive experiments demonstrated SemViQA’s state-of-the-art performance on ISE-DSC01 and ViWikiFC. Additionally, the SemViQA Faster variant accelerates inference by up to 7×, improving its practicality for real-world applications. By addressing key challenges such as semantic ambiguity and multi-step reasoning, SemViQA lays the groundwork for advancing Vietnamese NLP, with potential applications in misinformation detection and low-resource language fact-checking.
Limitations
While SemViQA demonstrates strong performance in Vietnamese fact verification, several limitations remain. First, our reliance on TF-IDF for initial evidence retrieval, while efficient, limits the model’s ability to capture deep semantic relationships and retrieve implicit evidence. To mitigate this, we employ a threshold-based mechanism to identify hard samples and process them with a more advanced retrieval model. However, this approach relies on manually defined thresholds, which may not generalize well across different datasets, underscoring the need for adaptive and data-driven retrieval strategies in future work. Second, our Two-step Verdict Classification (TVC) framework improves claim verification accuracy but requires multiple classification stages, increasing inference time compared to single-step approaches. This additional computational cost is particularly significant in three-class classification tasks, where optimizing model efficiency without compromising accuracy remains a key challenge. Future work should focus on refining retrieval mechanisms and classification strategies to enhance efficiency and robustness, ensuring broader applicability of SemViQA in real-world fact verification scenarios.
Appendix A Strict Accuracy in Fact-Checking
Strict Accuracy: This metric is a stringent measure that requires both the verdict and the evidence to be predicted correctly compared to the ground truth sample.
Verdict ($v$ and $v^{\prime}$): refers to the verdict of the sample and the predicted verdict (supported, refuted, nei).
Evidence ($e$ and $e^{\prime}$): refers to the evidence of the sample and the predicted evidence.
$StrAcc=f(v,v^{\prime}).f(e,e^{\prime})$ (11)
Where:
$f(v,v^{\prime})=\begin{cases}1&\text{if }v=v^{\prime}\\ 0&\text{otherwise}\end{cases}$ (12)
$f(e,e^{\prime})=\begin{cases}1&\text{if }e=e^{\prime}\\ 0&\text{otherwise}\end{cases}$ (13)
Strict accuracy is the average of all StrAcc values.
Appendix B Hyperparameter and LLM Training Configuration
In this section, we present the detailed hyperparameter settings and training configurations for both our SemViQA models and the Large Language Model (LLM) fine-tuning process. Table 4 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering Token Classifier (QATC), and LLM fine-tuning.
Hyperparameter BC TC QATC LLM
Epochs 20 20 20 1
RT Loss - - -
Cross-Entropy Loss - -
Focal Loss - - -
Learning Rate $1e^{-5}$ $1e^{-5}$ $2e^{-6}$ $5e^{-5}$
Batch Size 104 104 36 2
Gradient Accumulation 1 1 2 1
Optimizer (AdamW)
Max Token Length 256 256 512 4096
GPUs A100 A100 A100 4 × H100
Zero - - - Zero3
LR Schedule Linear Linear Cyclic Cosine
Mixed Precision - - - bf16
Table 4: Consolidated hyperparameter and training configuration for SemViQA models and LLM fine-tuning.
We fine-tune a Large Language Model (LLM) using a restructured version of the original datasets, ViWikiFC and ISE-DSC01, as detailed in Figure 5. These datasets have been carefully adapted for training to improve performance and ensure compatibility with our model. For training, we utilize the official Qwen LLM implementation from the QwenLM repository [^4] https://github.com/QwenLM/Qwen. Our training setup follows the full configuration outlined in Table 4, ensuring optimal efficiency and alignment with best practices.
{mdframed}
[backgroundcolor=gray!5, linewidth=0.75pt, roundcorner=5pt]
Question: You are tasked with verifying the correctness of the following statement.
- We provide you with a claim and a context. Please classify the claim into one of three labels: “SUPPORTED”, “REFUTED”, or “NEI” (Not Enough Info).
- Your answer should include the classification label and the most relevant evidence sentence from the context.
- Remember, the evidence must be a full sentence, not part of a sentence or less than one sentence.
Given a claim and context as follows:
Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day. For dinner, Ivy Chen usually eats early to ensure her body has enough time to digest food, metabolize energy, and avoid putting pressure on the stomach and other organs. A recent study published in *Frontiers in Nutrition* suggests that eating dinner earlier can lead to a longer lifespan, with the ideal time being 7 PM. If this is not possible, experts recommend having the last meal of the day 2-3 hours before bedtime. Drinking ginger tea: To keep her body warm, promote blood circulation, and enhance circulation, Ivy Chen drinks ginger tea daily. Her ginger tea is typically made with ground ginger, black tea, turmeric powder, and brown sugar. This drink is a natural remedy that not only boosts the immune system and reduces inflammation but also fights oxidation, supports weight loss, improves skin health, and helps maintain a youthful look. Regular exercise: Ivy Chen is a fitness enthusiast who loves physical activities and exercises daily, even during pregnancy. The Taiwanese actress shared that if she is not busy with work, she runs for at least 30 minutes every day. Even when traveling abroad, she maintains her running habit. A recent study published in *Progress in Cardiovascular Disease* found that regular runners live three years longer than non-runners. Running significantly helps with weight loss, maintaining a balanced physique, toning muscles, relaxing the mind, and benefiting heart health. Besides running, Ivy Chen also swims, practices yoga, and hikes to maintain physical fitness and endurance. Skincare: Regarding her skincare routine, the actress emphasized the importance of hydration. The Taiwanese beauty revealed that she always carries a facial mist to ensure her skin stays hydrated while outdoors.
Claim: Even when traveling abroad, Ivy Chen maintains her running habit.
Answer: This claim is classified as SUPPORTED. The evidence is: Even when traveling abroad, she maintains her running habit.
Table 5: Example of a fact-checking task prompt used for LLM training. Note: Some parts of the Context and Claim were originally in Vietnamese. In this paper, we have translated them into English for better readability. Sentences highlighted in blue indicate the evidence.
[…]
Figure 7: Training progress of the Qwen 1.5B and Qwen 3B models.
Appendix C Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection
Claim Evidence TF-IDF QATC
Du lịch Triều Tiên là điều mà chỉ có một số người được đi đến.
(Traveling to North Korea is something only a few people can do.)
Theo nguyên tắc, bất kỳ ai cũng được phép du lịch tới Triều Tiên, và những ai có thể hoàn thành quá trình làm thủ tục thì đều không bị Triều Tiên từ chối cho nhập cảnh.
(In principle, anyone is allowed to travel to North Korea, and those who complete the process are not denied entry.)
Khách du lịch không được đi thăm thú bên ngoài vùng đã được cho phép trước mà không được hướng dẫn viên người Triều Tiên cho phép nhằm tránh các điệp viên nằm vùng.
(Tourists are not allowed to visit areas outside of the designated zones without a North Korean guide to prevent undercover spies.)
Theo nguyên tắc, bất kỳ ai cũng được phép du lịch tới Triều Tiên, và những ai có thể hoàn thành quá trình làm thủ tục thì đều không bị Triều Tiên từ chối cho nhập cảnh.
(In principle, anyone is allowed to travel to North Korea, and those who complete the process are not denied entry.)
Nó có độ nóng chảy ở mức gần 30 độ C.
(It has a melting point of about 30°C.)
Nó là một kim loại kiềm mềm, màu bạc, và với điểm nóng chảy là 28 °C (83 °F) khiến cho nó trở thành một trong các kim loại ở dạng lỏng tại hay gần nhiệt độ phòng.
(It is a soft, silvery alkali metal with a melting point of 28°C (83°F), making it one of the metals that is liquid at or near room temperature.)
Nó là nguyên tố có độ âm điện thấp thứ hai sau franci, và chỉ có một đồng vị bền là caesi-133.
(It is the second least electronegative element after francium, and has only one stable isotope, cesium-133.)
Nó là một kim loại kiềm mềm, màu bạc, và với điểm nóng chảy là 28 °C (83 °F) khiến cho nó trở thành một trong các kim loại ở dạng lỏng tại hay gần nhiệt độ phòng.
(It is a soft, silvery alkali metal with a melting point of 28°C (83°F), making it one of the metals that is liquid at or near room temperature.)
Table 6: Comparison of TF-IDF and QATC in Fact-Checking: TF-IDF selects irrelevant evidence (Incorrect), while QATC selects accurate evidence (Correct).