### v3

# SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking

Abstract
Recent advances in LLMs have accelerated both information generation and misinformation, especially in low-resource languages like Vietnamese, motivating robust fact-checking systems. Existing methods struggle with semantic ambiguity, homonyms, and complex linguistic structures, often trading accuracy for efficiency. We introduce SemViQA, a novel Vietnamese fact-checking framework integrating Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC). Our approach balances precision and speed, achieving state-of-the-art results with 78.97% strict accuracy on ISE-DSC01 and 80.82% on ViWikiFC, securing 1st place in the UIT Data Science Challenge. Additionally, SemViQA Faster improves inference speed 7× while maintaining competitive accuracy. SemViQA sets a new benchmark for Vietnamese fact verification, advancing the fight against misinformation. The source code is available at: https://github.com/DAVID-NGUYEN-S16/SemViQA.
SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking
1 Introduction
The rapid advancement of large language models (LLMs), such as OpenAI’s ChatGPT, Google Gemini Team et al. (2024), Llama3.1 Touvron et al. (2023), Qwen2.5 Qwen et al. (2025), DeepSeek V3, DeepSeek-AI et al. (2024), Phi3.5 Abdin et al. (2024) has significantly improved information retrieval and processing across various domains. However, a major challenge with these systems is their tendency to generate factually incorrect or hallucinated content seemingly plausible information that lacks factual grounding Soleimani et al. (2020). This issue is particularly critical in domains requiring high accuracy, such as healthcare, law, and journalism, where misinformation can have serious consequences. Consequently, developing reliable fact-checking systems capable of retrieving and evaluating evidence from real-world sources has become an urgent need in Natural Language Processing (NLP).
Although fact verification has been widely studied in high-resource languages like English, applying these methods to low-resource languages such as Vietnamese remains a significant challenge. Transformer-based models, including BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), have demonstrated strong performance but their adaptation to Vietnamese is still limited. ViNSV Tran et al. (2024b) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but suffers from SBERT’s 256-token input constraint, making it ineffective for complex, long-context claims. Graph-based reasoning methods Zhong et al. (2020) offer promising semantic inference but are often computationally expensive. Traditional retrieval methods like TF-IDF and BM25, while efficient, rely heavily on exact keyword matching, limiting their ability to capture nuanced semantics. Recent large language model (LLM) approaches Huo et al. (2023); Schimanski et al. (2024) show potential but typically require substantial computational resources, creating a trade-off between speed and accuracy.
To address these challenges, we propose SemViQA, a Vietnamese fact-checking framework that balances semantic accuracy and computational efficiency. As shown in Figure 1, SemViQA comprises three key components:
1.
Semantic-based Evidence Retrieval (SER): Includes a preprocessing step that efficiently handles long-token contexts by splitting them into manageable subcontexts (e.g., 400 tokens). It combines fast TF-IDF retrieval with selective Question Answering Token Classification (QATC) to strike a balance between speed and semantic accuracy.
2.
Two-step Verdict Classification (TVC): Employs a hierarchical classification strategy with both three-class and binary classification stages to enhance robustness and improve performance on challenging claim verification tasks.
SemViQA achieves 78.97% strict accuracy on ISE-DSC01 111https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC Le et al. (2024), outperforming existing baselines (see Table 1). These results validate SemViQA’s potential to enhance Vietnamese fact verification, supporting misinformation mitigation and improved transparency.
The rest of this paper is organized as follows: Section 2 reviews related work, Section 3 presents the methodology, Section 4 reports experimental results, and Section 5 concludes the paper with future directions.
2 Related Works
Advances in Natural Language Processing (NLP) have driven rapid progress in fact verification and evidence extraction. Early BiLSTM‑based models such as the Neural Semantic Matching Network (NSMN) Nie et al. (2018)—augmented with WordNet features improved accuracy but struggled with complex sentence relations due to sequential limitations Graves and Schmidhuber (2005). Transformer models, notably BERT Devlin et al. (2019), introduced bidirectional contextual encoding and achieved state‑of‑the‑art results on FEVER Soleimani et al. (2020); Zhou et al. (2019); Malon (2018); Aly et al. (2021); Lin et al. (2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), yet their input‑length cap hinders long‑document fact‑checking. Graph‑based reasoning further enhances multi‑hop verification GCN variants Zhong et al. (2020); Thorne et al. (2018) and AdMIRaL’s logic‑driven retrieval Aly and Vlachos (2022) boost evidence sufficiency but at considerable computational cost. Vietnamese research is sparse: ViNSV Tran et al. (2024b) pairs BM25 with SBERT Reimers and Gurevych (2019a) yet falters on complex reasoning due to static embeddings.
TF-IDF remains the industry standard for document retrieval due to its speed, simplicity, interpretability, and ability to handle long contexts effectively, making it suitable for rapid and scalable retrieval Reddy et al. (2018); Qaiser and Ali (2018); Li (2021); Azevedo et al. (2022). However, its reliance on surface-level keyword matching limits its capability to handle paraphrases, contextual nuances, and multi-hop reasoning, reducing recall accuracy on complex queries. Ensemble learning Hannichenko et al. (2023); Wang et al. (2021); Liu et al. (2024); Ganaie et al. (2022) mitigates individual model weaknesses by aggregating diverse architectures and training signals, yielding robust gains. Building on these insights, SemViQA fuses fast TF‑IDF retrieval, semantic reasoning via QATC, and hierarchical classification, delivering high accuracy, low latency, and practical scalability for Vietnamese fact verification.
3 SemViQA - Semantic Vietnamese Question Answering
We formulate Vietnamese fact verification as a multi-output classification task, where the input is a pair , with being a claim and its corresponding context paragraph or document. The objective is to (i) identify the most relevant evidence sentence(s) from and (ii) predict the veracity label of the claim as one of three categories: Supported, Refuted, or Not Enough Information (NEI). To address challenges such as long input sequences and semantic ambiguity, we propose SemViQA, a three-stage framework consisting of data pre-processing, semantic-based evidence retrieval, and two-step verdict classification. An overview of the architecture is shown in Figure 2, with detailed descriptions provided in the following subsections.
3.1 Evidence Extraction via Question Answering with Token Classification
3.1.1 Data Processing
To effectively support evidence retrieval and claim classification, we apply distinct preprocessing strategies to the context based on the specific requirements of each downstream task (see Appendix A). A key challenge arises from the considerable length of many context passages, frequently exceeding the token limits of Vietnamese BERT-based models. Figure 3 illustrates our approach for handling long input contexts. First, the context is segmented into individual sentences. Next, sentences are sequentially aggregated into subcontexts until reaching approximately 400 tokens. Each completed subcontext is then processed by the QATC model to identify potential evidence. The next subcontext begins from the subsequent sentence, and this process continues until all sentences are processed. However, processing subcontexts sequentially can be time-consuming. Therefore, we developed SemViQA Faster, which batches and processes subcontexts in parallel, significantly accelerating the retrieval process.
3.1.2 Question Answering with Token Classification (QATC)
Traditional Question Answering models typically predict the start and end positions of an answer span. In our framework, we enhance this approach by introducing a token-level classification objective, enabling the model to focus explicitly on tokens within evidence sentences in the context. This dual formulation provides improved supervision for evidence extraction. Drawing inspiration from rationale tagging Ju et al. (2019), we treat token labeling as a binary classification task: tokens within evidence sentences receive a label of 1, and all other tokens receive a label of 0. In cases marked NEI (Not Enough Information), every token is labeled as 0. We employ a feed-forward classification layer on the token representations:
where is the contextual representation of token , are learnable weights in the neural network and denotes the sigmoid function. The loss for this task is the Binary Cross-Entropy (BCE) loss:
3.2 Semantic-based Evidence Retrieval (SER)
Accurate claim verification requires reliable evidence. To improve both efficiency and robustness, we adopt a two-stage evidence retrieval strategy combining TF-IDF with a QATC.
Stage 1: TF-IDF-based Retrieval. We segment the context into smaller passages and pair each with the claim . Preprocessing includes noise removal and tokenization using ViTokenizer222https://github.com/trungtv/pyvi. TF-IDF is effective for simple claims particularly refuted ones but struggles with semantically complex cases due to its reliance on keyword overlap. To enrich short segments (i.e., those with fewer than 60% of ’s tokens), we merge them with preceding segments to improve evidence completeness. Retrieved segments are then ranked, and a confidence threshold is applied to identify easy cases (handled by TF-IDF) and hard cases (passed to QATC).
Stage 2: QATC-based Refinement. For complex cases, QATC is applied to segmented subcontexts rather than the full context due to the input length limitation of BERT models. The detailed processing approach is described in Section 3.1.1. At this time, we consider three scenarios: (1) If multiple subcontexts yield conflicting answers, we collect all predicted spans and re-rank them using TF-IDF. (2) If a single evidence span consistently appears, it is selected directly. (3) If no evidence is found, fallback to TF-IDF is used.
This hybrid approach balances speed and semantic accuracy, improving evidence selection for downstream verdict classification. Examples are provided in Appendix E.
3.3 Two-step Verdict Classification (TVC)
We adopt a two-stage classification framework to enhance claim verification robustness and mitigate label imbalance, especially the overrepresentation of NEI.
Stage 1: Three-Class Classification.
Given a claim-evidence pair , a BERT-based model predicts a probability distribution over three labels: Supported, Refuted, and Not Enough Information (NEI):
This step is optimized using Cross-Entropy Loss.
Stage 2: Binary Classification.
If , we apply a refined binary classifier to distinguish between Supported and Refuted:
This model uses Focal Loss Lin et al. (2018) to address class imbalance.
Final Prediction Rule.
The final verdict where is determined by comparing the confidence scores from both classifiers. Here, represents the index of the predicted label, where each index corresponds to a specific class description.
where , represents the highest probability.
This hybrid strategy allows the three-class model to handle general cases, especially detecting NEI early, while the binary model specializes in distinguishing difficult SUP/REF cases.
3.4 SemViQA Pipeline System
We now describe the full SemViQA pipeline, as illustrated in Figure 2. First, we prepare input for TF-IDF by splitting the context paragraph into sentences, then concatenating each sentence with the claim . We calculate the matching score for each sentence and select the one with the highest probability. If this score exceeds the threshold , we directly use this sentence as the evidence. If the score is below , we proceed to prepare input for the QATC model. We segment into subcontexts (as detailed in Section 3.1.1). Each subcontext is sequentially processed by the QATC model. If QATC identifies zero or multiple candidate evidence spans, we collect all predicted spans and re-rank them using TF-IDF. If QATC finds exactly one candidate evidence span, we confidently select it as the final evidence. Finally, we move to the two-step verdict classification (TVC) stage. We prepare input for TVC by concatenating the claim with the final evidence. We first apply the three-class model. If it predicts NEI, the process ends. Otherwise, we use an ensemble method to combine the weights from both the three-class and binary models to make the final prediction. All coefficients used in our system are provided in the ablation study, as shown in Appendix C.
4 Experiments
4.1 Dataset
We use two Vietnamese fact verification datasets: ISE-DSC01 from the UIT Challenge 2023 and ViWikiFC Le et al. (2024), which contains over 20,000 Wikipedia-based claims, including annotated evidence for the “nei” label. Dataset details are provided in Appendix A.1.
| Method | ViWikiFC | ISE-DSC01 | Avg Strict Acc | |||||||
| ER | VC | Strict Acc | VC Acc | ER Acc | Time (s) | Strict Acc | VC Acc | ER Acc | Time (s) | |
| Traditional Baselines | ||||||||||
| TF-IDF | InfoXLM | 75.56 | 82.21 | 90.15 | 131 | 73.59 | 78.08 | 76.61 | 378 | 74.58 |
| XLM-R | 76.47 | 82.78 | 90.15 | 134 | 75.61 | 80.50 | 78.58 | 366 | 76.04 | |
| Ernie-M | 75.56 | 81.83 | 90.15 | 144 | 78.19 | 81.69 | 80.65 | 403 | 76.88 | |
| BM25 | InfoXLM | 70.44 | 79.01 | 83.50 | 130 | 72.09 | 77.37 | 75.04 | 320 | 71.27 |
| XLM-R | 70.97 | 78.91 | 83.50 | 132 | 73.94 | 79.37 | 76.95 | 333 | 72.46 | |
| Ernie-M | 70.21 | 78.29 | 83.50 | 141 | 76.58 | 80.76 | 79.02 | 381 | 73.40 | |
| SBert | InfoXLM | 74.99 | 81.59 | 89.72 | 195 | 71.20 | 76.59 | 74.15 | 915 | 73.10 |
| XLM-R | 75.80 | 82.35 | 89.72 | 194 | 72.85 | 78.78 | 75.89 | 835 | 74.33 | |
| Ernie-M | 75.13 | 81.44 | 89.72 | 203 | 75.46 | 79.89 | 77.91 | 920 | 75.30 | |
| QA-based Approaches | ||||||||||
| ViMRC | InfoXLM | 77.28 | 81.97 | 92.49 | 3778 | 54.36 | 64.14 | 56.84 | 9798 | 65.82 |
| XLM-R | 78.29 | 82.83 | 92.49 | 3824 | 53.98 | 66.70 | 57.77 | 9809 | 66.14 | |
| Ernie-M | 77.38 | 81.92 | 92.49 | 3785 | 56.62 | 62.19 | 58.91 | 9833 | 67.00 | |
| InfoXLM | InfoXLM | 78.14 | 82.07 | 93.45 | 4092 | 53.50 | 63.83 | 56.17 | 10057 | 65.82 |
| XLM-R | 79.20 | 83.07 | 93.45 | 4096 | 53.32 | 66.70 | 57.25 | 10066 | 66.26 | |
| Ernie-M | 78.24 | 82.21 | 93.45 | 4102 | 56.34 | 62.36 | 58.69 | 10078 | 67.29 | |
| LLMs | ||||||||||
| Qwen2.5-1.5-Instruct | 51.03 | 65.18 | 78.96 | 7665 | 59.23 | 66.68 | 65.51 | 19780 | 55.13 | |
| Qwen2.5-3B-Instruct | 44.38 | 62.31 | 71.35 | 12123 | 60.87 | 66.92 | 66.10 | 31284 | 52.63 | |
| Qwen2.5-1.5-Instruct | InfoXLM | 66.14 | 76.47 | 78.96 | 7788 | 64.40 | 68.37 | 66.49 | 19970 | 65.27 |
| XLM-R | 67.67 | 78.10 | 78.96 | 7789 | 64.66 | 69.63 | 66.72 | 19976 | 66.17 | |
| Ernie-M | 66.52 | 76.52 | 78.96 | 7794 | 65.70 | 68.37 | 67.33 | 20003 | 66.11 | |
| Qwen2.5-3B-Instruct | InfoXLM | 59.88 | 72.50 | 71.35 | 12246 | 65.72 | 69.66 | 67.51 | 31477 | 62.80 |
| XLM-R | 60.74 | 73.08 | 71.35 | 12246 | 66.12 | 70.44 | 67.83 | 31483 | 63.43 | |
| Ernie-M | 60.02 | 72.21 | 71.35 | 12251 | 67.48 | 70.77 | 68.75 | 31512 | 63.80 | |
| Ours: SER Faster + TVC | ||||||||||
| ViMRC | Ernie-M | 79.44 | 82.93 | 94.60 | 410 | 78.32 | 81.91 | 80.26 | 995 | 78.88 |
| InfoXLM | 79.77 | 83.07 | 95.03 | 487 | 78.37 | 81.91 | 80.32 | 925 | 79.07 | |
| Ours: Full SER + TVC | ||||||||||
| ViMRC | InfoXLM | 80.25 | 83.84 | 94.69 | 2731 | 75.13 | 79.54 | 76.87 | 5191 | 77.69 |
| XLM-R | 80.34 | 83.64 | 94.69 | 2733 | 76.71 | 81.65 | 78.91 | 5219 | 78.53 | |
| Ernie-M | 79.53 | 82.97 | 94.69 | 2733 | 78.97 | 82.54 | 80.91 | 5225 | 79.25 | |
| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 3860 | 75.13 | 79.60 | 76.87 | 5175 | 77.91 |
| XLM-R | 80.82 | 83.88 | 95.31 | 3843 | 76.74 | 81.71 | 78.95 | 5200 | 78.78 | |
| Ernie-M | 80.06 | 83.17 | 95.31 | 3891 | 78.97 | 82.49 | 80.91 | 5297 | 79.52 | |
4.2 Experimental Setup
We conducted extensive experiments on NVIDIA A100 GPUs, fine-tuning key hyperparameters while keeping consistent settings across runs. The final configuration, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC. Full details are provided in Appendix D. For fair evaluation, all methods were tested on a Kaggle instance with an NVIDIA T4 GPU.
The large language model was fine-tuned in a distributed A100 setup using a structured prompt-based reformulation. Raw data were converted into prompt format to align with LLM training objectives and maximize task-specific performance. Training setup, prompt design, and preprocessing are also detailed in Appendix D.
4.3 Main Results
The results in Table 1 demonstrate that SemViQA outperforms previous methods in Vietnamese fact-checking tasks. Specifically, our model achieves the highest Strict Accuracy, reaching 80.82% on ViWikiFC and 78.97% on ISE-DSC01, establishing a new benchmark for automated fact-checking systems in Vietnamese language.
4.3.1 Performance Comparison
a) Handling Long Token Sequences in Fact-Checking
A major limitation of conventional Question Answering (QA) models in fact verification is their inability to process long-context claims due to the 512-token input limit of transformer-based models such as ViMRClarge333https://huggingface.co/nguyenvulebinh/vi-mrc-large, InfoXLMlarge Chi et al. (2021), XLM-Rlarge Conneau et al. (2020), and Ernie-Mlarge Ouyang et al. (2021). Real-world datasets like ISE-DSC01 often contain contexts exceeding 4800 tokens, severely degrading QA-based performance by limiting access to full evidence. To overcome this, SemViQA employs an efficient retrieval-based strategy (see Section 3.1.1) that handles long-token sequences effectively. On ISE-DSC01, SemViQA outperforms traditional QA models by fully leveraging extended contexts, confirming that the long-token constraint is a critical bottleneck. Conversely, on ViWikiFC, where contexts average around 512 tokens, QA models perform competitively. Yet, even in this setting, integrating our Semantic-based Evidence Retrieval (SER) yields a 1.86% improvement in evidence retrieval accuracy, demonstrating the versatility and efficiency of our approach. These findings emphasize that long-token limitations significantly hinder fact verification, and SemViQA successfully mitigates this issue while enhancing QA models across varying dataset conditions.
b) Performance and Inference Time Optimization
SemViQA significantly reduces inference time while maintaining high accuracy, making it highly practical for real-world applications. Key highlights include:
On ISE-DSC01, SemViQA averages 5200s per run, over 6 times faster than large LLM-based models like Qwen2.5-3B-Instruct Qwen et al. (2024), which require over 31,000s.
Compared to ViMRClarge (9800s), SemViQA halves inference time while achieving superior Strict Accuracy and Veracity Classification Accuracy.
Although BM25 and SBERT Reimers and Gurevych (2019b) are faster, they struggle with complex, multi-step reasoning where SemViQA maintains a strong balance between speed and accuracy.
SemViQA Faster: We further introduce SemViQA Faster, which accelerates inference by up to 7 through batch processing of subcontexts (see Section 3.1.1). As shown in Figure 4, this variant achieves inference speeds comparable to traditional methods while retaining competitive accuracy. The minor performance trade-off is acceptable given the substantial time savings, making SemViQA Faster ideal for scalable, real-world fact-checking systems.
4.3.2 Comparison with other results in the competition
| Methods | Strict Acc | VC Acc | ER Acc |
| SemViQA | 78.97 | 82.54 | 80.91 |
| DS@UIT Dynasty444https://github.com/minhquan6203/Verdict-Classification-for-Fact-Checking-at-DSC-2023 | 78.05 | 84.76 | 80.13 |
| URA_FNU555https://github.com/virrosluo/URA_UIT_Data_Science_Challenge | 77.87 | 83.71 | 79.96 |
| ViNSV Tran et al. (2024b) | 76.33 | 81.67 | 78.11 |
| Tran et al. (2024a) | 75.11 | 82.30 | 76.82 |
The results presented in Table 2 indicate that our SemViQA approach outperforms other competing teams, achieving the highest Strict Accuracy and demonstrating exceptional effectiveness in information processing and verification. This achievement highlights SemViQA’s capability to deliver significantly more accurate and reliable results.
5 Conclusion and Future Works
We introduced SemViQA, a Vietnamese fact-checking framework that integrates Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC) to enhance claim verification. Our approach outperforms existing methods, including LLMs, TF-IDF, BM25, SBERT, and QA-based models, particularly in handling long-token sequences and complex reasoning tasks. Extensive experiments demonstrated SemViQA’s state-of-the-art performance on ISE-DSC01 and ViWikiFC. Additionally, the SemViQA Faster variant accelerates inference by up to 7×, improving its practicality for real-world applications. By addressing key challenges such as semantic ambiguity and multi-step reasoning, SemViQA lays the groundwork for advancing Vietnamese NLP, with potential applications in misinformation detection and low-resource language fact-checking.
Limitations
While SemViQA demonstrates strong performance in Vietnamese fact verification, several limitations remain. First, our reliance on TF-IDF for initial evidence retrieval, while efficient, limits the model’s ability to capture deep semantic relationships and retrieve implicit evidence. To mitigate this, we employ a threshold-based mechanism to identify hard samples and process them with a more advanced retrieval model. However, this approach relies on manually defined thresholds, which may not generalize well across different datasets, underscoring the need for adaptive and data-driven retrieval strategies in future work. Second, our Two-step Verdict Classification (TVC) framework improves claim verification accuracy but requires multiple classification stages, increasing inference time compared to single-step approaches. This additional computational cost is particularly significant in three-class classification tasks, where optimizing model efficiency without compromising accuracy remains a key challenge. Future work should focus on refining retrieval mechanisms and classification strategies to enhance efficiency and robustness, ensuring broader applicability of SemViQA in real-world fact verification scenarios.
Appendix A Comprehensive Data Processing
A.1 Dataset Statistics
This section presents the detailed statistics of the two datasets used in our experiments: ISE-DSC01 and ViWikiFC. The table below summarizes the number of samples in the training, development, and test sets for both corpora. These statistics help contextualize the scale and evaluation coverage of our methods.
| ISE-DSC01 | ViWikiFC | |
| Train | 37,967 | 16,738 |
| Dev | 4,794 | 2,090 |
| Test | 5,396 | 2,091 |
A.2 Data Analysis on Context Lengths
Figure 5 illustrates the distribution of token lengths across input contexts in the ViWikiFC and ISE-DSC01 datasets. As shown, many samples significantly exceed the 512-token input limitation of standard Transformer models. The ISE-DSC01 dataset, in particular, contains several contexts with over 4,000 tokens. This analysis highlights the necessity for effective context segmentation strategies to ensure full coverage of relevant evidence while maintaining compatibility with model constraints.
Figure 3 illustrates our strategy for handling long input contexts. First, the context is segmented into individual sentences. Next, sentences are sequentially aggregated into subcontexts until reaching approximately 400 tokens. Each completed subcontext is then passed to the QATC model to identify potential evidence. The next subcontext begins from the subsequent sentence, continuing until all sentences have been processed. However, processing subcontexts sequentially can be time-consuming. Therefore, we developed SER Faster, which batches subcontexts and processes them in parallel, significantly accelerating the retrieval process.
Appendix B Strict Accuracy in Fact-Checking
Strict Accuracy: This metric is a stringent measure that requires both the verdict and the evidence to be predicted correctly compared to the ground truth sample.
Verdict ( and ): refers to the verdict of the sample and the predicted verdict (supported, refuted, nei).
Evidence ( and ): refers to the evidence of the sample and the predicted evidence.
Where:
Strict accuracy is the average of all StrAcc values.
Appendix C Ablation Study
C.1 Ablation Experiments
| Configuration | ViWikiFC | ISE-DSC01 | |||||
| ER Model | VC Model | Strict Acc | VC Acc | ER Acc | Strict Acc | VC Acc | ER Acc |
| SemViQA (Full) | |||||||
| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 75.13 | 79.60 | 76.87 |
| XLM-R | 80.82 | 83.88 | 95.31 | 76.74 | 81.71 | 78.95 | |
| Ernie-M | 80.06 | 83.17 | 95.31 | 78.97 | 82.49 | 80.91 | |
| w/o Binary Classification (One-step VC) | |||||||
| InfoXLM | InfoXLM | 79.63 | 82.88 | 95.31 | 73.87 | 78.35 | 76.89 |
| XLM-R | 80.73 | 83.69 | 95.31 | 75.96 | 80.80 | 78.97 | |
| Ernie-M | 79.91 | 83.07 | 95.31 | 78.47 | 81.89 | 80.93 | |
| w/o QATC (TF-IDF-based ER) | |||||||
| TF-IDF | InfoXLM | 76.57 | 83.26 | 90.15 | 74.89 | 79.36 | 76.61 |
| XLM-R | 76.47 | 82.93 | 90.15 | 76.39 | 81.41 | 78.58 | |
| Ernie-M | 75.75 | 81.97 | 90.15 | 78.71 | 82.28 | 80.65 | |
Table 4 presents the ablation results to evaluate the contribution of each component in the SemViQA framework. When employing the full model with both the QATC-based evidence retrieval and the two-step verdict classification (TVC), SemViQA achieves the best performance across both datasets. Notably, using Ernie-Mlarge yields the highest strict accuracy (78.97%) and evidence retrieval accuracy (80.91%) on the ISE-DSC01 dataset. Removing the binary classification stage (i.e., using only one-step VC) leads to a noticeable performance drop, especially on ISE-DSC01, indicating that the binary classifier enhances the distinction between Supported and Refuted labels. Furthermore, replacing the QATC module with a TF-IDF based retriever results in a significant decline (about 5%) in evidence retrieval accuracy, which subsequently affects the overall performance. These findings highlight the critical role of both QATC and the two-step classification scheme in improving SemViQA’s effectiveness on fact verification in Vietnamese.
C.2 Analysis of Confidence Threshold in SemViQA
The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process. Analysis from Figure 6 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%). However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively. Setting an optimal threshold in the range of 0.4 - 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
C.3 Effectiveness of QATC over Traditional QA-based Evidence Retrieval
| SemViQA | ViWikiFC | ISE-DSC01 | |||||
| SER | TVC | Strict Acc | VC Acc | ER Acc | Strict Acc | VC Acc | ER Acc |
| QATC-based ER | |||||||
| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 75.13 | 79.60 | 76.87 |
| XLM-R | 80.82 | 83.88 | 95.31 | 76.74 | 81.71 | 78.95 | |
| Ernie-M | 80.06 | 83.17 | 95.31 | 78.97 | 82.49 | 80.91 | |
| ViMRClarge | InfoXLMlarge | 80.25 | 83.84 | 94.69 | 75.13 | 79.54 | 76.87 |
| XLM-Rlarge | 80.34 | 83.64 | 94.69 | 76.71 | 81.65 | 78.91 | |
| Ernie-Mlarge | 79.53 | 82.97 | 94.69 | 78.97 | 82.54 | 80.91 | |
| QA-based ER | |||||||
| InfoXLM | InfoXLM | 79.96 | 83.50 | 94.45 | 74.02 | 78.95 | 75.83 |
| XLM-R | 80.11 | 83.60 | 94.45 | 75.61 | 80.95 | 77.91 | |
| Ernie-M | 79.24 | 82.74 | 94.45 | 77.82 | 81.76 | 79.82 | |
| ViMRClarge | InfoXLMlarge | 79.77 | 83.84 | 94.26 | 74.05 | 78.93 | 75.87 |
| XLM-R | 79.87 | 83.79 | 94.26 | 75.65 | 80.93 | 77.95 | |
| Ernie-M | 79.01 | 82.78 | 94.26 | 77.84 | 81.73 | 79.86 | |
To compare the learning capabilities of QATC with traditional QA models, we construct two versions of the SemViQA pipeline that are identical in structure, differing only in the evidence retrieval model. As shown in Table 5, using QATC consistently outperforms QA-based retrieval across both datasets. Specifically, on ViWikiFC, QATC achieves up to 80.82% Strict Accuracy and 95.31% ER Accuracy, while QA-based models peak at 80.11% and 94.45%, respectively. The improvement is even more evident on ISE-DSC01, where QATC reaches 78.97% Strict Accuracy and 80.91% ER Accuracy. These results confirm that QATC is more effective at learning to identify relevant evidence, leading to superior performance across the entire fact-checking pipeline.
C.4 Analysis of Confusion Matrix in Verdict Classification
To evaluate the effectiveness of our Two-step Verdict Classification (TVC) strategy, we analyze the confusion matrices in Figure 7, which compare the standard three-class classification with our proposed two-step approach, focusing solely on Supported and Refuted claims. The three-class classifier achieves strong performance, with accuracy of 95.4% for Supported and 94.6% for Refuted; however, it still exhibits notable confusion between the two classes, likely due to their semantic proximity and shared evidence patterns. In contrast, the two-step approach—by isolating NEI cases early and applying a dedicated binary classifier further improves accuracy to 98.0% for Supported and 96.4% for Refuted. These results support our hypothesis that decomposing the verdict classification task into semantically coherent subtasks enhances the model’s precision in detecting factual consistency.
Appendix D Hyperparameter and LLM Training Configuration
In this section, we present the detailed hyperparameter settings and training configurations for both our SemViQA models and the Large Language Model (LLM) fine-tuning process. Table 6 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering with Token Classification (QATC), and LLM fine-tuning.
| Hyperparameter | BC | TC | QATC | LLM |
| Epochs | 20 | 20 | 20 | 1 |
| RT Loss | - | - | ✓ | - |
| Cross-Entropy Loss | - | ✓ | ✓ | - |
| Focal Loss | ✓ | - | - | - |
| Learning Rate | ||||
| Batch Size | 104 | 104 | 36 | 2 |
| Gradient Accumulation | 1 | 1 | 2 | 1 |
| Optimizer (AdamW) | ✓ | ✓ | ✓ | ✓ |
| Max Token Length | 256 | 256 | 512 | 4096 |
| GPUs | A100 | A100 | A100 | A100 |
| Zero | - | - | - | Zero3 |
| LR Schedule | Linear | Linear | Cyclic | Cosine |
| Mixed Precision | - | - | - | bf16 |
We fine-tune a Large Language Model (LLM) using a restructured version of the original datasets, ViWikiFC and ISE-DSC01, as detailed in Figure 7. These datasets have been carefully adapted for training to improve performance and ensure compatibility with our model. For training, we utilize the official Qwen LLM implementation from the QwenLM repository666https://github.com/QwenLM/Qwen. Our training setup follows the full configuration outlined in Table 6, ensuring optimal efficiency and alignment with best practices.
- We provide you with a claim and a context. Please classify the claim into one of three labels: “Supported”, “REFUTED”, or “NEI” (Not Enough Info).
- Your answer should include the classification label and the most relevant evidence sentence from the context.
- Remember, the evidence must be a full sentence, not part of a sentence or less than one sentence. Given a claim and context as follows: Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day. For dinner, Ivy Chen usually eats early to ensure her body has enough time to digest food, metabolize energy, and avoid putting pressure on the stomach and other organs. A recent study published in *Frontiers in Nutrition* suggests that eating dinner earlier can lead to a longer lifespan, with the ideal time being 7 PM. If this is not possible, experts recommend having the last meal of the day 2-3 hours before bedtime. Drinking ginger tea: To keep her body warm, promote blood circulation, and enhance circulation, Ivy Chen drinks ginger tea daily. Her ginger tea is typically made with ground ginger, black tea, turmeric powder, and brown sugar. This drink is a natural remedy that not only boosts the immune system and reduces inflammation but also fights oxidation, supports weight loss, improves skin health, and helps maintain a youthful look. Regular exercise: Ivy Chen is a fitness enthusiast who loves physical activities and exercises daily, even during pregnancy. The Taiwanese actress shared that if she is not busy with work, she runs for at least 30 minutes every day. Even when traveling abroad, she maintains her running habit. A recent study published in *Progress in Cardiovascular Disease* found that regular runners live three years longer than non-runners. Running significantly helps with weight loss, maintaining a balanced physique, toning muscles, relaxing the mind, and benefiting heart health. Besides running, Ivy Chen also swims, practices yoga, and hikes to maintain physical fitness and endurance. Skincare: Regarding her skincare routine, the actress emphasized the importance of hydration. The Taiwanese beauty revealed that she always carries a facial mist to ensure her skin stays hydrated while outdoors. Claim: Even when traveling abroad, Ivy Chen maintains her running habit. Answer: This claim is classified as Supported. The evidence is: Even when traveling abroad, she maintains her running habit.
We present the complete training progress of the LLM models and QATC in Figure 9 and Figure 8, respectively. Figure 9 illustrates the training dynamics of Qwen 1.5B and Qwen 3B, supporting the results presented in Table 1. Notably, the Qwen 1.5B model demonstrates more stable training dynamics compared to the Qwen 3B model during the initial stage. Meanwhile, Figure 8 showcases the completion of QATC training, depicting the loss curves of ViMRC and InfoXLM. These results highlight the convergence behavior of QATC training across different architectures, further supporting the robustness of our approach.
Appendix E Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection
| Claim | Evidence | TF-IDF | QATC |
Du lịch Triều Tiên là điều mà chỉ có một số người được đi đến.
(Traveling to North Korea is something only a few people can do.) |
Theo nguyên tắc, bất kỳ ai cũng được phép du lịch tới Triều Tiên, và những ai có thể hoàn thành quá trình làm thủ tục thì đều không bị Triều Tiên từ chối cho nhập cảnh.
(In principle, anyone is allowed to travel to North Korea, and those who complete the process are not denied entry.) |
Khách du lịch không được đi thăm thú bên ngoài vùng đã được cho phép trước mà không được hướng dẫn viên người Triều Tiên cho phép nhằm tránh các điệp viên nằm vùng.
(Tourists are not allowed to visit areas outside of the designated zones without a North Korean guide to prevent undercover spies.) |
Theo nguyên tắc, bất kỳ ai cũng được phép du lịch tới Triều Tiên, và những ai có thể hoàn thành quá trình làm thủ tục thì đều không bị Triều Tiên từ chối cho nhập cảnh.
(In principle, anyone is allowed to travel to North Korea, and those who complete the process are not denied entry.) |
Nó có độ nóng chảy ở mức gần 30 độ C.
(It has a melting point of about 30°C.) |
Nó là một kim loại kiềm mềm, màu bạc, và với điểm nóng chảy là 28 °C (83 °F) khiến cho nó trở thành một trong các kim loại ở dạng lỏng tại hay gần nhiệt độ phòng.
(It is a soft, silvery alkali metal with a melting point of 28°C (83°F), making it one of the metals that is liquid at or near room temperature.) |
Nó là nguyên tố có độ âm điện thấp thứ hai sau franci, và chỉ có một đồng vị bền là caesi-133.
(It is the second least electronegative element after francium, and has only one stable isotope, cesium-133.) |
Nó là một kim loại kiềm mềm, màu bạc, và với điểm nóng chảy là 28 °C (83 °F) khiến cho nó trở thành một trong các kim loại ở dạng lỏng tại hay gần nhiệt độ phòng.
(It is a soft, silvery alkali metal with a melting point of 28°C (83°F), making it one of the metals that is liquid at or near room temperature.) |